OpenStudy (anonymous):

|12|+|-11| 1/2+.5 1/8+.125 |-4+-3| |-18+3|

OpenStudy (anonymous):

23 1 .25 7 15

OpenStudy (anonymous):

The x/x's represent fractions.

OpenStudy (anonymous):

yes - but i gave answers as decimals - do you want them to be as fractions?

OpenStudy (anonymous):

Yes, please.

OpenStudy (anonymous):

ok second = 1/2 + 1/2 = 1 3rd = 1/8 + 1/8 = 1/4 (.5 = 1/2 and .125 = 1/8)

OpenStudy (anonymous):

Cheers, mate.

OpenStudy (anonymous):

no probs